OpenStudy (anonymous):

Express 5t in terms of the base 2 exponential function

OpenStudy (anonymous):

\[f(t)=5^t\] like that?

OpenStudy (anonymous):

and you want to use the base \(2\) rather than \(5\) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

only thing i can think of is to solve \[5^x=2\] for \(x\)

OpenStudy (anonymous):

by the change of base formula you get \[x=\frac{\ln(2)}{\ln(5)}\] making \[\large 5^{\frac{\ln(2)}{\ln(5)}}=2\] and so \[\left(\large 5^{\frac{\ln(2)}{\ln(5)}}\right)^t=2^t\]

OpenStudy (anonymous):

oh but maybe that is backwards!

OpenStudy (anonymous):

yes, it is backwards i think solve \[2^x=5\] so \[x=\frac{\ln(5)}{\ln(2)}\]etc etc

OpenStudy (anonymous):

was thinkinking the same, i am pretty confused...i hate it when math problems are worded, but thanks

OpenStudy (anonymous):

yeah, it is confusing but i think what you want is \[\left(\large 2^{\frac{\ln(5)}{\ln(2)}}\right)^t=5^t\]